OpenStudy (anonymous):

Will fan and medal! Please help! An inspector inspects large loads of potatoes to determine the proportion p in the shipment with major defects prior to making potato chips. She selects an SRS of 50 potatoes from over 2000 in the load. Only 2 of the potatoes sampled are found to have major defects. Suppose you instead wish to test these hypotheses: H0:mu=0.10 HAmu<0.10 What is the p-value for this test? Choose one answer. A. 0.9835 B. 0.1573 C. 0.0786 D. 0.9214 E. zero

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Are they giving you a table to find the p value?

OpenStudy (anonymous):

No, that question is all the information I was given.

jimthompson5910 (jim_thompson5910):

Do they expect you to use a table in the back of your book maybe?

OpenStudy (anonymous):

I am taking the course online thus there is no back of the book.

jimthompson5910 (jim_thompson5910):

hmm frustrating

jimthompson5910 (jim_thompson5910):

they should have at least provided a table or a calculator maybe

jimthompson5910 (jim_thompson5910):

anyways, do you know how to find the z-score?

OpenStudy (anonymous):

Yes, but not with the information supplied.

jimthompson5910 (jim_thompson5910):

You use the formula z = (p - P)/(sqrt(P*(1-P)/n)) In this case p = 2/50 = 0.04 P = 0.10 (based on the null hypothesis) n = 50

OpenStudy (anonymous):

so z=-1.4142 ?

jimthompson5910 (jim_thompson5910):

now use this calculator http://www.danielsoper.com/statcalc3/calc.aspx?id=53

jimthompson5910 (jim_thompson5910):

in this case, mu = 0 sigma = 1

jimthompson5910 (jim_thompson5910):

then you type -1.4142 into that last box

OpenStudy (anonymous):

I got 0.07865159

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

that's the approximate p value

jimthompson5910 (jim_thompson5910):

basically the area under the curve to the left of -1.4142

OpenStudy (anonymous):

Great, thanks! Can you help with this one? A psychic was tested for ESP. The psychic was presented with 400 cards face down and was asked to determine if the card was one of four symbols: a cross, a star, a circle, or a square. The psychic was correct in 120 of the cases. Let p represent the probability that the psychic correctly identifies the symbol on the card in a random trial. How large a sample n would you need to estimate p with a margin of error of 0.01 with 95% confidence? (Use the guess 0.25 as the value of p) Choose one answer. A. N = 447 B. N = 7203 C. N = 9604 D. N = 30

jimthompson5910 (jim_thompson5910):

check out this page http://www.ltcconline.net/greenl/courses/201/estimation/ciprop.htm

jimthompson5910 (jim_thompson5910):

what you need is at the very bottom, but it doesn't hurt to read the other stuff (since you may need it someday)

OpenStudy (anonymous):

What do zsubc and E represent?

jimthompson5910 (jim_thompson5910):

E is the margin of error

jimthompson5910 (jim_thompson5910):

zsubc is the z critical value

jimthompson5910 (jim_thompson5910):

in the case of the 95% confidence interval, it's roughly 1.96 since 95% of the distribution is between z = -1.96 and z = 1.96

OpenStudy (anonymous):

So how do I get the critical value?

jimthompson5910 (jim_thompson5910):

Well you either * memorize it * use a table (an inverse cdf table) * use a calculator

jimthompson5910 (jim_thompson5910):

here's a table you can use http://3.bp.blogspot.com/_5u1UHojRiJk/TEdJJc6of2I/AAAAAAAAAIE/Ai0MW5VgIhg/s1600/t-table.jpg

jimthompson5910 (jim_thompson5910):

notice in the row that starts with \(\Large \infty\), we have 1.960 right above the 95% confidence interval

OpenStudy (anonymous):

So is 1.960 the critical value?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

So I went through it and got n=7203, is that correct?

jimthompson5910 (jim_thompson5910):

I got the same 0.25*0.75*(1.96/0.01)^2 = 7,203